OpenStudy (anonymous):

Solve 6x + 10 – 2x = 4(x + 2).

OpenStudy (anonymous):

6x+10-2x=4x+8 6x-4x=2x 2x+10-2x=+8 2x+2x=4x 4x+10=8 8-10=-2 4x=-2 -2/4=-0.5